OpenStudy (anonymous):

Carla has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 10 inches long, and 7 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? The paint will not fill the tray by 44.59 in3. The paint will not fill the tray by 469 cm3. The paint will fill the tray exactly. The paint will overfill the tray by 44.59 in3.

OpenStudy (phi):

can you find the volume of the tray?

OpenStudy (phi):

using V= L*W*H

OpenStudy (anonymous):

1/2.54=x/7 (7cm to inches) x=7/2.54 V(tray)=10*10*(7/2.54) V(tray)= 35000/127 = 275.5905512 in3 1 in3 = 48391/2953 (16.387064) cm3 1 gallon=3785.41178 cm3 V(tray)= 4516.120001 cm3

OpenStudy (anonymous):

FLVS? You would first need to convert all measurements to the same unit. I would change them to inches. They give you the conversion 1 in = 2.54 cm. So divide 7 by 2.54 and that will give you how many inches 7 cm is.

OpenStudy (anonymous):

that is 2.8 now what

OpenStudy (anonymous):

To get the volume. you multiply width * length * height Multiply 10 * 10 * (7/2.54).

OpenStudy (anonymous):

We know that 1 gallon is 231 in^3. So now take the number you just got and subtract 231.

OpenStudy (anonymous):

okay so 280-231=49

OpenStudy (anonymous):

Sorta. It would be more accurate if you use 7/2.54 instead of 2.8. So 10*10*(7/2.54) = 275.59 Now subtract 231

OpenStudy (anonymous):

44.59

OpenStudy (anonymous):

yes thats what im doing now